OpenStudy (anonymous):

f(X) = (sqrt(x-3))/((x+8)(x-2) find domain of function

OpenStudy (campbell_st):

well looking at the denominator x cannot be -8 or 2 either value would result in a zero denominator which is undefined. now looking at the numerator x must be greater than or equal to 3 because if x is less than 3 you are looking at negative square roots. so the domain of the function is \[x \ge3\]

OpenStudy (anonymous):

see thats exactly what i thought, but these are my options x > 0 All real numbers x ≥ -3, x ≠ 2 x ≠ -8, x ≠ -3, x ≠ 2

OpenStudy (campbell_st):

are you sure the numerator is correct... looking at the answers its appears that the numerator is \[\sqrt{x + 3}\]

OpenStudy (anonymous):

sorry you are right, it is what + 3

OpenStudy (campbell_st):

so then what value can't you have in the numerator its \[x \ge -3\] combine this with the values in the denominator and you'll get a domain

OpenStudy (anonymous):

the closest one to that is the third answer but it doesnt have x isn't equal to -8

OpenStudy (campbell_st):

well x can't go as low as -8 so you are looking at \[x \ge -3,...x \ne2\]

OpenStudy (anonymous):

oh okay now i understand. thnakyou